# Comparison results for Markov tree distributions

### v1

@@ Appendix C Proofs of Section 2
    Similar to the proof of [80, Theorem 2], it follows that $\mu_{\mathcal{T}}$ is the unique distribution with the given specifications that has Markov tree dependence. For $|N|=\infty$, the distribution $\mu_{\mathcal{T}}$ can be obtained by the Kolmogorov extension Theorem, see [66, Theorem 6.16], where the assumption of projective families is satisfied as a consequence of the disintegration theorem. $\blacksquare$
Appendix D Proofs of Section 3
For a random vector $X=(X_{0},\dots,X_{d})$ with distribution function $F_{X}$, we denote by $\overline{F}_{X}$ the survival function of $X$ which is given by
$\overline{F}_{X}(x_{0},\dots,x_{d}):=P(X_{0}>x_{0},\dots,X_{d}>x_{d}).$
There is also an analogous version of Sklar’s Theorem for survival functions, cf. [42, Theorem 2.2.13], which states that for any survival function $\overline{F}_{X}$ with survival marginals $\overline{F}_{X_{n}}(x)\coloneqq 1-F_{X_{n}}(x)$, $x\in\mathbb{R}$, there is a copula $\widehat{C}_{X}$, such that
$\overline{F}_{X}(x_{0},\dots,x_{d})=\widehat{C}_{X}(\overline{F}_{X_{0}}(x_{0}),\dots,\overline{F}_{X_{d}}(x_{d})).$ (84)
The copula $\widehat{C}_{X}$ is called survival copula of $X$. The following lemma shows that the copula (survival copula) of a Markov tree distribution with respect to a tree $T=(N,E)$, specified by continuous marginal distribution functions $F_{i},\,i\in N,$ and bivariate copulas $B_{e},\,e\in E$, is given by the distribution function (the survival copula) of the Markov tree distribution specified by uniform marginals on $[0,1]$ and the bivariate copulas $B_{e},\,e\in E$. This is especially useful for comparing bivariate tree specification with the same bivariate copula structure but different marginal distributions.
Lemma D.1
@@ Appendix D Proofs of Section 3
Let $X=(X_{0},\ldots,X_{d})\sim\mathcal{M}((F_{0},\dots,F_{d}),T,C)$ be a random vector with continuous marginal specifications $F_{0},\dots,F_{d}$. Moreover, let $U=(U_{0},\ldots,U_{d})\sim\mathcal{M}((G,\dots,G),T,C)$ be a random vector with distribution function $F_{U}$, where $G$ denotes the distribution function of the uniform distribution on $[0,1]$. Then the following hold true:
  1. (i)
    $F_{U}$ is the unique copula for $X$, i.e., $F_{X}(x_{1},\dots,x_{d})=F_{U}(F_{1}(x_{1}),\dots,F_{d}(x_{d}))$.
  2. (ii)
    The survival function and the survival copula of $U$ are related by
    $\widehat{\mathcal{C}}_{U}(u_{0},\dots,u_{d})=\overline{F}_{U}(1-u_{0},\dots,1-u_{d}).$ (85)
    Further $\widehat{\mathcal{C}}_{U}$ is the unique survival copula of $X$, i.e., $\widehat{\mathcal{C}}_{U}$ is the unique copula such that
    $\widehat{\mathcal{C}}_{U}(\overline{F_{0}}(x_{0}),\dots,\overline{F_{d}}(x_{d}))=P(X_{0}>x_{0},\dots,X_{d}>x_{d}),$ (86)
    where $\overline{F}_{n}=\overline{F}_{X_{n}}$ is the survival function of $X_{n}\,,$ $n\in\{0,\ldots,n\}\,.$
  • Proof of Lemma D.1
@@
    We prove statement (i) by an induction over the number of nodes $N=\{0,\dots,d\}$. Starting with $|N|=2$, the statement is trivial. For the induction step, assume that the statement is true for each tree $T$ with nodes $N=\{0,\dots,n\}$ for a fixed $n\in\{1,\ldots,d-1\}\,.$ We need to show the statement for $N=\{0,\dots,n+1\}$. By re-enumeration we may assume that $n+1\in N$ is a leaf in $N$ with parent node $n$, i.e., $c_{n+1}=\emptyset$ and $p_{n+1}=n$, see Definition B.1. We define the subtree $T^{\prime}=(N^{\prime},E^{\prime})$ with nodes $N^{\prime}=\{0,\dots,n\}$ and edges $E^{\prime}=\{(j,k)\in E|j,k\neq n+1\}$ which is a tree with number of nodes $|N^{\prime}|-1=n$. By (83) and the transformation formula, using that $F_{i}^{-1}\circ F_{i}=\operatorname{id}$ $P^{X_{i}}$-a.s., cf. [101, Theorem 1.2], we have that
    $\begin{split}&F_{(X_{0},\dots,X_{n+1})}(x_{0},\dots,x_{n})\\ &=\int_{(-\infty,x_{0}]\times\dots\times(-\infty,x_{n}]}F_{X_{n+1}|X_{n}=y_{n}}(x_{n+1})P^{(X_{0},\dots,X_{n})}(dy_{0},\dots,dy_{n})\\ &=\int_{(-\infty,F_{0}(x_{0})]\times\dots\times(-\infty,F(x_{n})]}F_{X_{n+1}|X_{n}=F_{n}^{-1}(u_{n})}(x_{n+1})P^{(F_{0}(X_{0}),\dots,F_{n}(X_{n}))}(du_{0},\dots,du_{n}).\end{split}$ (87)
@@
    Since the marginal distribution functions are continuous, the distribution function of the transformed random vector $(F_{0}(X_{0}),\dots,F_{n}(X_{n}))$ is the unique copula of $(X_{0},\dots,X_{n})$, see, e.g., [99, Remark 2.2]. Moreover, by the induction hypothesis the distribution function $F_{U_{0},\dots,U_{n}}$ is also a copula for $(X_{0},\dots,X_{n})$. By uniqueness of the copula, we obtain that $(F_{0}(X_{0}),\dots,F_{n}(X_{n}))\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(U_{0},\dots,U_{n})$. By [6, Theorem 2.2], we have
    $F_{X_{n+1}|X_{n}=F_{n}^{-1}(u_{n})}(x_{n+1})=\partial_{1}C_{n,n+1}(u_{n},F_{n+1}(x_{n+1}))=F_{U_{n+1}|U_{n}=u_{n}}(F_{n+1}(x_{n+1})),$
    where $\partial_{1}$ denotes the operator that takes the partial derivative with respect to the first component. Hence, we obtain that (87) equals
@@
    $=F_{(U_{0},\dots,U_{n+1})}(F_{1}(x_{1}),\dots,F_{d}(x_{n+1})),$
    which proves the first result. We continue with statement (ii). Equation (85) follows directly by
    $\overline{F}_{U}(1-\overline{F}_{U_{0}}(u_{0}),\dots,1-\overline{F}_{U_{d}}(u_{d}))=\overline{F}_{U}(u_{0},\dots,u_{d}).$ (88)
    By [42, Theorem 2.2.13] and the continuity of $F_{0},\dots,F_{d}$ the survival copula for $X$ is unique and it remains to show equation (86). By the inclusion-exclusion principle we obtain that
    $\widehat{\mathcal{C}}_{U}(u_{0},\dots,u_{d})=1-\sum_{k=1}^{d+1}\left((-1)^{k+1}\sum_{I\subseteq N,|I|=k}F_{U_{I}}\Big{(}(1-u_{i})_{i\in I}\Big{)}\right),$ (89)
    see also [42, Equation 1.7.5]. Moreover, by the first statement (i) we obtain that
    $\begin{split}P(X_{0}>x_{0},\dots,X_{d}>x_{d})&=1-P\left(\bigcup_{n=0}^{d}\{X_{n}\leq x_{n}\}\right)\\ &=1-\sum_{k=1}^{d+1}\left((-1)^{k+1}\sum_{I\subseteq N,|I|=k}F_{X_{I}}\Big{(}(x_{i})_{i\in I}\Big{)}\right)\\ &=1-\sum_{k=1}^{d+1}\left((-1)^{k+1}\sum_{I\subseteq N,|I|=k}F_{U_{I}}\Big{(}(1-\overline{F}_{i}(x_{i}))_{i\in I}\Big{)}\right).\end{split}$ (90)
    Finally, (89) and (90) shows (86) and the second statement is proven. $\blacksquare$
  • Proof of Theorem 3.6
@@
    By the definition of the orders for stochastic processes, Definition 2.7 (b), we may assume that $|N|<\infty$. Let $Z=(Z_{n})_{n\in N}\sim\mathcal{M}(F,T,C)$ be a random vector having the same marginals as $X$ and the same bivariate dependence specifications as $Y\,.$ By assumption (ii) and (14) the copulas $C_{e}$, $e\in E,$ are CI and thus $(Z_{i},Z_{j})$ is CI for all $(i,j)\in E$. We deduce from Theorem 1.3 that $X\leq_{sm}Z$ and consequently $X\leq_{dcx}Z$. When verifying the assumptions for Theorem 1.3, it is worth noting that distinguishing between the cases where $\operatorname{deg}(0)\geq 2$ and $\operatorname{deg}(0)<2$ is not necessary since $(Y_{i},Y_{j})$ is CI for all $(i,j)\in E$ and consequently in the case where $\operatorname{deg}(0)\geq 2$, it is always possible to select the path $P$ in the assumptions of Theorem 1.3 in such a way that $k^{*}\notin P.$ According to [44, Proposition 7.1], the vectors $Z$ and $Y$ are MTP2 and thus CI. By the continuity of $F_{n}$ and $G_{n}$ for all $n\in N$ and Lemma D.1, the vectors $Z$ and $Y$ have the same copula which is CI. Hence, we obtain from [85, Theorem 4.5] using that $F_{n}\leq_{cx}G_{n}$ for all $n\in N$, that $Z\leq_{dcx}Y$. The transitivity of the directionally convex order implies $X\leq_{dcx}Y$. $\blacksquare$
  • Proof of Theorem 3.7
    Consider a sequence $Z=(Z_{n})_{n\in N}\sim\mathcal{M}(F,T,C)$ of Markov tree distributed random variables. Since $Y$ fulfills assumption (ii) of Theorem 1.3 this is also the case for $Z$. Moreover $X$ and $Z$ have the same marginal distributions. Hence, $X$ and $Z$ fulfill the assumptions of Theorem 1.3 and we obtain that $X\leq_{sm}Z$. Due to (8), it holds that $X\leq_{lo}Z$ and $X\leq_{uo}Z$. By the transitivity of the lower and upper orthant order it remains to show that $Z\leq_{lo}Y$ (resp. $Z\leq_{uo}Y$). Since the marginal distribution functions $F_{n}$ and $G_{n}$ are assumed to be continuous and since $Z$ and $Y$ have the same bivariate specifications, Lemma D.1 (i) implies that $Z$ and $Y$ have the same copula $\mathcal{C}\in\mathcal{C}_{d+1}$. Since, by assumption, $F_{n}(x)\leq G_{n}(x)$ for all $x\in\mathbb{R}\,,$ we obtain from Sklar’s Theorem that
    $F_{Z}(x)=\mathcal{C}(F_{0}(x_{0}),\dots,F_{d}(x_{d}))\leq\mathcal{C}(G_{0}(x_{0}),\dots,G_{d}(x_{d}))=F_{Y}(x)$
    for all $x=(x_{0},\ldots,x_{d})\in\mathbb{R}^{d+1}\,.$ This proves $Z\leq_{lo}Y\,.$ To show $Z\leq_{uo}Y\,,$ let $\widehat{\mathcal{C}}\in\mathcal{C}$ be the common survival copula of $Z$ and $Y$, see Lemma D.1 (ii). Since, by assumption, $\overline{F}_{n}(x)\leq\overline{G}_{n}(x)$ for all $x\in\mathbb{R}\,,$ we obtain
    $\overline{F}_{Z}(x)=\widehat{\mathcal{C}}(\overline{F}_{0}(x_{0}),\dots,\overline{F}_{d}(x_{d}))\leq\widehat{\mathcal{C}}(\overline{G}_{0}(x_{0}),\dots,\overline{G}_{d}(x_{d}))=\overline{F}_{Y}(x)$
    for all $x=(x_{0},\ldots,x_{d})\in\mathbb{R}^{d+1}\,,$ which concludes the proof. $\blacksquare$
Appendix E Proofs of Section 4
  • Proof of Corollary 4.1
    Hidden Markov models follow a Markov tree distribution with respect to the tree $T=(\mathbb{N}_{0},E)$, where $E$ is defined by Equation (17). The first statement then follows by Theorem 1.3. The second and third statement are a direct consequence of Theorem 3.7. The fourth statement follows from Theorem 3.6. $\blacksquare$
Appendix F Proofs of Section 5
  • Proof of Proposition 5.2
(a) For the first statement, we consider two cases. If $j^{*}\in d_{k^{*}}$, i.e., $j^{*}$ is an element of the decedents of $k^{*}$, we define the vectors by $(X_{0},X_{k^{*}},X_{p(k^{*},j^{*})},X_{j^{*}})\coloneqq-Y^{\prime}$ and $(Y_{0},Y_{k^{*}},Y_{p(k^{*},j^{*})},Y_{j^{*}})\coloneqq-X^{\prime}$, where $X^{\prime}$ and $Y^{\prime}$ are given by Example A.5. Then we extend these vectors by independent random variables (note that the independence copula is CI) to the Markov tree distributed sequences $X=(X_{i})_{i\in N}$ and $Y=(Y_{i})_{i\in N}$. These vectors fulfill assumption ($i^{*}$) in (33), and the assumptions (ii) and (iii) of Theorem 1.3. Since $(X_{0},X_{k^{*}},X_{p(k^{*},j^{*})},X_{j^{*}})\not\leq_{lo}(Y_{0},Y_{k^{*}},Y_{p(k^{*},j^{*})},Y_{j^{*}})$, we obtain from the closure of the lower orthant order under marginalization that $X\nleq_{lo}Y$ and thus $X\nleq_{sm}Y\,.$ In the case where $j^{*}\notin d_{k^{*}}$, which implies $\operatorname{deg}(0)\geq 2$, we define the vectors $(X_{k^{*}},X_{0},X_{p(0,j^{*})},X_{j^{*}})\coloneqq-Y^{\prime}$ and $(Y_{k^{*}},Y_{0},Y_{p(0,j^{*})},Y_{j^{*}})\coloneqq-X^{\prime}$, and proceed as in the first case.
(b) Since $T$ is not a star it holds $N\setminus(L\cup\{0\})\neq\emptyset$. The node $j^{*}\in N\setminus(L\cup\{0\})\,,$ has a parent node $i\coloneqq p_{j^{*}}$ and has at least one child node $k\in c_{j^{*}}$. Define the subvectors $(X_{i},X_{j^{*}},X_{k})$ and $(Y_{i},Y_{j^{*}},Y_{k})$ as in Example A.3 and extend these $3$-dimensional vectors by independent random variables (note that the independence copula is CI) to the random vectors $X=(X_{i})_{i\in N}$ and $Y=(Y_{i})_{i\in N}$. Then $X$ and $Y$ fulfill assumption ($ii^{*}$) in (34), and the assumptions (i) and (iii) of Theorem 1.3. We obtain from Example A.3 that $(X_{i},X_{j^{*}},X_{k})\nleq_{lo}(Y_{i},Y_{j^{*}},Y_{k})$ and thus $X\nleq_{lo}Y$ and $X\nleq_{sm}Y\,$. This shows the second statement.
(c) Since $T$ is not a chain, we have $N\backslash(P\cup\{0\})\neq\emptyset$. For $j^{*}\in N\backslash(P\cup\{0\})$ we consider the directed path $(0,\dots,j^{*})$ from the root 00 to $j^{*}$. Let $i$ be the last node in $(0,\dots,j^{*})$ such that $i\in P\cup\{0\}$, i.e., the unique node $i\in\{0,\dots,j^{*}\}\cap(P\cup\{0\})$ such that $c_{i}\cap\{0,\dots,j^{*}\}\cap P=\emptyset$. Since $i\in P$ and $j^{*}\notin P$ the node $i$ has a child node $k\in P\cap c_{i}$. Define the subvectors $(X_{k},X_{p(k,j^{*})},X_{j^{*}})$ and $(Y_{k},Y_{p(k,j^{*})},Y_{j^{*}})$ as in Example A.5 and extend this $3$-dimensional vectors by independent random variables (note that the independence copula is CI) to the random vectors $X=(X_{i})_{i\in N}$ and $Y=(Y_{i})_{i\in N}$. Then $X$ and $Y$ fulfill assumption ($ii^{**}$) in (35), and the assumptions (i) and (iii) of Theorem 1.3. Due to Example A.5, we obtain $(X_{k},X_{p(k,j^{*})},X_{j^{*}})\nleq_{lo}(Y_{k},Y_{p(k,j^{*})},Y_{j^{*}})$ and consequently $X\nleq_{lo}Y$ and $X\nleq_{sm}Y\,$. This shows the third statement and the result is proven. $\blacksquare$
  • Proof of Proposition 5.4

### v3

@@ B.2 Proofs of Section 2
Similar to the proof of [54, Theorem 2], it follows that $\mu_{\mathcal{T}}$ is the unique distribution with the given specifications that has Markov tree dependence. For $|N|=\infty$, the distribution $\mu_{\mathcal{T}}$ can be obtained by the Kolmogorov extension Theorem (see [45, Theorem 6.16]), where the assumption of projective families is satisfied as a consequence of the disintegration theorem. $\blacksquare$
B.3 Proofs of Section 3
For the proofs of Theorems 3.1 and 3.5, we make use of the following lemma.
Lemma B.6
@@ B.3 Proofs of Section 3
Let $X=(X_{0},\ldots,X_{d})\sim\mathcal{M}((F_{0},\dots,F_{d}),T,C)$ be a random vector with marginal specifications $F_{0},\dots,F_{d}$. Further, let $U=(U_{0},\ldots,U_{d})\sim\mathcal{M}((G_{0},\dots,G_{d}),T,C)$ be a random vector with distribution function $F_{U}$, where $G_{i}$ denotes the distribution function of the rank-transformed random variable $F_{i}(X_{i}).$ Then we have
$F_{X}(x_{0},\dots,x_{d})=F_{U}(F_{0}(x_{0}),\dots,F_{d}(x_{d}))$ (81)
for all $(x_{0},\ldots,x_{d})\in\mathbb{R}^{d+1}.$
@@
We prove the statement by an induction over the number of nodes $N=\{0,\dots,d\}$. Starting with $|N|=2$, the statement is trivial. For the induction step, assume that the statement is true for each tree $T$ with nodes $N=\{0,\dots,n\}$ for a fixed $n\in\{1,\ldots,d-1\}.$ We need to show the statement for $N=\{0,\dots,n+1\}$. By re-enumeration we may assume that $n+1\in N$ is a leaf in $N$ with parent node $n$, i.e., $c_{n+1}=\emptyset$ and $p_{n+1}=n$; see Definition B.1. We define the subtree $T^{\prime}=(N^{\prime},E^{\prime})$ with nodes $N^{\prime}=\{0,\dots,n\}$ and edges $E^{\prime}=\{(j,k)\in E|j,k\neq n+1\}$ which is a tree with number of nodes $|N^{\prime}|-1=n$. By (80) and the transformation formula, using that $F_{i}^{-1}\circ F_{i}(X_{i})=X_{i}$ $P$-a.s. (see e.g. [5, Lemma A.1(ii)]), we have that
$\begin{split}&F_{(X_{0},\dots,X_{n+1})}(x_{0},\dots,x_{n},x_{n+1})\\ &=\int_{(-\infty,x_{0}]\times\dots\times(-\infty,x_{n}]}F_{X_{n+1}|X_{n}=y_{n}}(x_{n+1})P^{(X_{0},\dots,X_{n})}(dy_{0},\dots,dy_{n})\\ &=\int_{(-\infty,F_{0}(x_{0})]\times\dots\times(-\infty,F(x_{n})]}F_{X_{n+1}|X_{n}=F_{n}^{-1}(u_{n})}(x_{n+1})P^{(F_{0}(X_{0}),\dots,F_{n}(X_{n}))}(\mathsf{\,d}\mathbb{u})\end{split}$ (82)
@@
for $\mathsf{\,d}\mathbb{u}=(du_{0},\dots,du_{n}).$ Since copulas are invariant under componentwise increasing transformations, it follows that the random vectors $(F_{0}(X_{0}),\dots,F_{n}(X_{n}))$ and $(X_{0},\dots,X_{n})$ have the same copula (see e.g. [17, Theorem 3.3]). Moreover, by the induction hypothesis the distribution function $F_{U_{0},\dots,U_{n}}$ coincides with the copula of $(X_{0},\dots,X_{n})$ on $\mathsf{Ran}(F_{0})\times\cdots\times\mathsf{Ran}(F_{d}).$ By uniqueness of the copula on $\mathsf{Ran}(F_{0})\times\cdots\times\mathsf{Ran}(F_{d}),$ we obtain that $(F_{0}(X_{0}),\dots,F_{n}(X_{n}))\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(U_{0},\dots,U_{n})$. Then we have
$F_{X_{n+1}|X_{n}=F_{n}^{-1}(u_{n})}(x_{n+1})$ $=\partial_{1}^{F_{n}}C_{n,n+1}(F_{n}(F_{n}^{-1}(u_{n})),F_{n+1}(x_{n+1}))$
$=\partial_{1}^{G_{n}}C_{n,n+1}(G_{n}(G_{n}^{-1}(u_{n})),F_{n+1}(x_{n+1}))$
$=\partial_{1}^{G_{n}}C_{n,n+1}(G_{n}(G_{n}^{-1}(u_{n})),G_{n+1}(F_{n+1}(x_{n+1})))$
$=F_{U_{n+1}|U_{n}=u_{n}}(F_{n+1}(x_{n+1})),$
for Lebesgue-almost all $u_{n}\in\mathsf{Ran}(F_{n})$ and for all $x_{n+1}\in\mathsf{Ran}(F_{n+1}),$ where the first and last equality hold true by [5, Theorem 2.2]. Here $\partial_{1}^{H}$ denotes a generalized partial derivative operator that takes the partial derivative with respect to the first component and with respect to a univariate distribution function $H,$ see [5, Section 2.1]. For the second equality, we use the assumptions that $(X_{n},X_{n+1})$ and $(U_{n},U_{n+1})$ have the same copula, and that $\overline{\mathsf{Ran}(F_{n})}=\overline{\mathsf{Ran}(G_{n})}$, noting that $\partial_{1}^{H}$ has the property that it only depend on $\overline{\mathsf{Ran}(H)}.$ For the third equality, we use that $G_{n+1}(v)=v$ for all $v\in\mathsf{Ran}(G_{n+1}).$ Hence, we obtain that (82) equals
@@
$=F_{(U_{0},\dots,U_{n+1})}(F_{0}(x_{0}),\dots,F_{d}(x_{n+1}))$
for all $(x_{0},\ldots,x_{n+1})\in\mathbb{R}^{n+2},$ which proves the statement. $\blacksquare$
By the definition of the orderings for stochastic processes, Definition 2.5 (b), we may assume that $|N|<\infty$. Let $Z=(Z_{n})_{n\in N}\sim\mathcal{M}(F,T,C).$ Then we have $X\leq_{sm}Z$ due to Theorem 1.3 noting that, due to (6), $B_{e}\leq_{lo}C_{e},$ $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z_{i},$ and $X_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z_{j}$ implies $(X_{i},X_{j})\leq_{sm}(Z_{i},Z_{j})$ for $e=(i,j).$ Now, observe that $Z$ and $Y$ have the same bivariate copula specifications. Since $\overline{\mathsf{Ran}(F_{n})}=\overline{\mathsf{Ran}(G_{n})}$ and thus $F_{n}(Z_{n})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}G_{n}(Y_{n})$ for all $n$, we obtain from Lemma B.6 that $Z$ and $Y$ have a common copula. Using the assumption that $Z_{n}\leq_{st}Y_{n}$ (resp. $Z_{n}\geq_{st}Y_{n}$) for all $n,$ it follows that $Z\leq_{st}Y$ (resp. $Z\geq_{st}Y$) see [59, Theorem 4.1]. Hence, for $f$ increasing (decreasing) supermodular, we obtain $\mathbb{E}f(X)\leq\mathbb{E}f(Z)\leq\mathbb{E}f(Y)$ whenever the expectations exist. $\blacksquare$
@@
Again, we may assume that $|N|<\infty$. Let $Z=(Z_{n})_{n\in N}\sim\mathcal{M}(F,T,C)$ be a random vector having the same univariate marginals as $X$ and the same bivariate dependence specifications as $Y.$ By assumption (ii) and (8) the copulas $C_{e}$, $e\in E,$ are CI and thus $(Z_{i},Z_{j})$ is CI for all $(i,j)\in E$. We deduce from Theorem 1.3 that $X\leq_{sm}Z$ and consequently $X\leq_{dcx}Z$. When verifying the assumptions for Theorem 1.3, it is worth noting that distinguishing between the cases where $\operatorname{deg}(0)\geq 2$ and $\operatorname{deg}(0)<2$ is not necessary since $(Y_{i},Y_{j})$ is CI for all $(i,j)\in E$. Consequently, in the case where $\operatorname{deg}(0)\geq 2$, it is always possible to select the path $P$ in the assumptions of Theorem 1.3 in such a way that $k^{*}\notin P.$ By the continuity of $F_{n}$ and $G_{n}$ for all $n\in N$ and Lemma B.6, the vectors $Z$ and $Y$ have the same copula. Using the MTP2 assumption, we obtain from Lemma 3.4, that this copula is MTP2, and thus CI. Hence, it follows with Lemma 3.3, using the assumption $F_{n}\leq_{cx}G_{n}$ for all $n\in N$, that $Z\leq_{dcx}Y$. The transitivity of the directionally convex order implies $X\leq_{dcx}Y$. $\blacksquare$
B.4 Proofs of Section 4